### 12 Jan 2024 v1

# WisdoM[Uncaptioned image]: Improving Multimodal Sentiment Analysis by Fusing Contextual World Knowledge

Abstract
Sentiment analysis is rapidly advancing by utilizing various data modalities (e.g., text, image). However, most previous works relied on superficial information, neglecting the incorporation of contextual world knowledge (e.g., background information derived from but beyond the given image and text pairs) and thereby restricting their ability to achieve better multimodal sentiment analysis. In this paper, we proposed a plug-in framework named WisdoM[Uncaptioned image] , designed to leverage contextual world knowledge induced from the large vision-language models (LVLMs) for enhanced multimodal sentiment analysis. WisdoM[Uncaptioned image] utilizes a LVLM to comprehensively analyze both images and corresponding sentences, simultaneously generating pertinent context. To reduce the noise in the context, we also introduce a training-free Contextual Fusion mechanism. Experimental results across diverse granularities of multimodal sentiment analysis tasks consistently demonstrate that our approach has substantial improvements (brings an average +1.89 F1 score among five advanced methods) over several state-of-the-art methods. Code will be released.
WisdoM[Uncaptioned image]: Improving Multimodal Sentiment Analysis by Fusing Contextual World Knowledge
Wenbin Wang1 Liang Ding2 Li Shen3 Yong Luo1 Han Hu4 Dacheng Tao2 1Wuhan University 2The University of Sydney 3JD Explore Academy 4Beijing Institute of Technology {wangwenbin97, luoyong}@whu.edu.cn, liangding.liam@gmail.com
1 Introduction
Sentiment analysis (SA, Medhat et al., 2014; Wankhade et al., 2022) focuses on identifying human sentiment polarity Das and Singh (2023). With the development of the Internet, the way people express their sentiments is not limited to only text, but also includes multimodal data (e.g., images). It is hard to detect the sentiment due to the short and informal nature of the sentence sometimes, but the associated image may help reflect people’s sentiment. How to use the information between different modalities becomes the main problem of multimodal sentiment analysis (MSA).
To address this problem, the relevant modal fusion strategies are carefully designed for MSA tasks of different granularity. For aspect-level MSA tasks, existing methods estimate whether the image contains information conducive to multimodal sentiment analysis by learning the relationships between image and text Ju et al. (2021); Ling et al. (2022); Yang et al. (2022); Zhou et al. (2023). For sentence-level MSA tasks, existing methods primary focus lies in the fusion at the feature level and decision level Hazarika et al. (2020); Yu et al. (2021); Han et al. (2021); Guo et al. (2022)
Figure 1: The simple schematic of our method. The sentiment polarity of Aleppo is negative, which is hard to directly predict by existing methods while our WisdoMRefer to caption predict correctly via incorporating context generated by the world knowledge-rich LVLMs.
Despite empirically success of above fusion techniques, there lacks investigations about improving MSA with (easy-to-acquired) world knowledge. This raises the following question: Could world knowledge boost MSA? For instance, in Fig. 1, given a comparative image at different periods alongside a sentence. Question: What’s the sentiment polarity of “Aleppo”? Current state-of-the-art (SOTA) multimodal sentiment analysis model (MSAM) Zhou et al. (2023) can only discern that the image is a comparative of Aleppo at various period. The predicted sentiment polarity is neutral. This is because the model lacks the knowledge of the Aleppo. In fact, human would associate Aleppo with being a city in Syria, and in conjunction with the image, we might infer connections to the Syrian war. Consequently, the deduction would be that the intent is to express the severe devastation experienced by the city of Aleppo due to the Syrian war, highlighting a stark contrast between its once-prosperous city and its current dilapidation. Hence, the sentiment towards Aleppo is negative.
Inspired by the human knowledge-guided in learning, we propose a plug-in framework WisdoM[Uncaptioned image] , utilizing contextual world knowledge from large vision-language models for MSA. Specifically, the WisdoM[Uncaptioned image] contains three stages: ❶Prompt Templates Generation, ❷Context Generation, and ❸Contextual Fusion. In stage 1, we ask ChatGPT [^1] https://chat.openai.com to generate prompt templates which are used to construct instructions for stage 2. The goal in stage 2 is to enables the large vision-language model (LVLM) Liu et al. (2023b) to generate contextual information based on the provided image and sentence. Note that, we refer to this contextual information as context. However, the context were found to be noisy in MSA. We further introduce a training-free Contextual Fusion that incorporating context for hard sample in stage 3.
We validated our WisdoM[Uncaptioned image] on several benchmarks including Twitter2015, Twitter2017 Yu and Jiang (2019) and MSED Jia et al. (2022) over several models: LLaVA-v1.5 Liu et al. (2023a), MMICL Zhao et al. (2023), mPLUG-Owl2 Ye et al. (2023), AoM Zhou et al. (2023), ALMT Zhang et al. (2023). Experiments demonstrate the effectiveness and universality of our approach. Extensive analysis confirms that WisdoM[Uncaptioned image] effectively enhances the ability of MSAM on MSA tasks.
Contributions
Our main contributions are:
  • We propose a plug-in framework WisdoM[Uncaptioned image] , leveraging the LVLM to generate explicit contextual world knowledge, to wisely enhance the multimodal sentiment analysis ability.
  • To achieve wise knowledge fusion, we introduce a novel contextual fusion mechanism to mitigate the impact of noise in the context.
  • We conduct experiments on three multimodal sentiment analysis benchmarks upon several advanced VLP models, showing that WisdoM[Uncaptioned image] could bring consistent and significant improvements (up to +5.8 F1 score).
2 Related Work
2.1 Multimodal Sentiment Analysis
Different from traditional text-based sentiment analysis task Hussein (2018), multimodal sentiment analysis (MSA) task Soleymani et al. (2017) integrates more modalities simultaneously, e.g., image and speech, to obtain accurate sentiment labels. Numerous advanced models have been proposed, covering different levels of granularity, such as sentence and aspect Jia et al. (2022). The core of most MSA studies lies in developing better fusion algorithms across modalities.
Sentence-level MSA
Wang et al. (2014) introduce a cross-media bag-of-words approach in the joint image-text framework for microblog images. You et al. (2015) employ deep neural networks to conduct a combined textual and visual sentiment analysis. Zhao et al. (2019) present a sentiment analysis model driven by image-text consistency, exploring correlations between image and text data in a movie review dataset. Li et al. (2020) propose a ConvTransformer, a multi-window convolutional transformer model that combines Transformer Vaswani et al. (2017) and CNN to identify joint sentiment polarity. Das and Singh (2022) propose a multi-stage multimodal sentiment analysis model for the Assamese language, utilizing a deep learning architecture that concurrently exploits textual and visual cues to determine sentiment. Zhang et al. (2023) propose an Adaptive Language-guided Multimodal Transformer (ALMT), which incorporates an Adaptive Hyper-modality Learning (AHL) module to learn an irrelevance/conflict-suppressing representation from different modalities under the guidance of language features at different scales.
Aspect-level MSA
Multimodal aspect-oriented sentiment classification combines both multimodal sentiment analysis Bezemer and Jewitt (2010) and aspect-oriented sentiment classification Wang et al. (2022); Zhong et al. (2023a), typically applied to multimedia content sourced from online social media Khan and Fu (2021). Yu and Jiang (2019) introduce TomBERT, leveraging two multimodal tweet datasets with target annotations, and proposed an aspect-oriented multimodal BERT model. TomBERT extends the baseline BERT architecture Kenton and Toutanova (2019) by incorporating target-sensitive visual attention and augmenting self-attention layers to capture inter-modality dynamics. Khan and Fu (2021) propose a two-stream model employing an object-aware transformer for image translation in the input space, followed by a single-pass non-autoregressive generation approach Wu et al. (2020); Ding et al. (2021).
Although existing approaches relying on sophisticated fusion techniques have achieved remarkable performance in MSA, their limitation lies in relying superficial information for fusion, without incorporating contextual world knowledge.
2.2 Large Vision-Language Models
Large Vision-Language Models (LVLMs) are becoming a fundamental tool for general task solver Li et al. (2023); Fu et al. (2023); Liu et al. (2023b); Zhao et al. (2023); Ye et al. (2023); Dai et al. (2023); Zhu et al. (2023); Alayrac et al. (2022); Chen et al. (2023). Liu et al. (2023a) introduce LLaVA, which establishes a connection between the pre-trained CLIP ViT-L/14 visual encoder Dosovitskiy et al. (2020) and the large language model Vicuna Chiang et al. (2023) through a straightforward projection matrix. The model is trained using a two-stage instruction-tuning procedure. Addressing the challenge of complex multi-modal prompts with multiple images, Zhao et al. (2023) propose MMICL, considering both model and data perspectives to enhance the understanding of LVLMs. Ye et al. (2023) propose mPUG-Owl2, a versatile multi-modal large language model that effectively utilizes modality collaboration to enhance performance in both text and multi-modal tasks. mPLUG-Owl2 employs a modularized network design, with the language decoder serving as a universal interface for managing various modalities.
Here, we explore the utility of existing LVLMs in generating contextual world knowledge for imporving multimodal sentiment analysis. Additionally, we introduce a training-free module, Contextual Fusion, to reduce the noise in the context.
Figure 2: Template of task instruction.
Figure 3: Detailed illustration of our proposed schema WisdoMRefer to caption with a running example. ❶ Using ChatGPT to provide prompt templates. ❷ We then prompt LVLMs to generate context using the prompt templates with image and sentence. ❸ A training-free mechanism Contextual Fusion mitigate the noise in the context.
3 Preliminary
In this section, we first describe the notation of the MSA. We then review two typical frameworks to modeling the MSA tasks, where we experiment our schema upon them: task-specific framework Zhou et al. (2023); Ling et al. (2022); Ju et al. (2021) and general-purpose framework Li et al. (2023).
Notation
Let $\mathcal{M}$ be a set of multimodal samples. Each sample $m_{i}\in\mathcal{M}$ consists of a sentence $s_{i}$ and image $v_{i}$. For aspect-level MSA tasks, there are several aspects $a_{i}$ which is a subsequence of $s_{i}$, i.e., $a_{i}\in s_{i}$. We denote $f(\cdot)$ as the sentiment classifier. The output of $f(\cdot)$ is the sentiment polarity $y_{i}\in$negative, neutral, positive}, with corresponding predicted probability denoted as $P_{i}=\{p^{neg}_{i},p^{neu}_{i},p^{pos}_{i}\}$.
Task-Specific Framework
For aspect-level MSA tasks, the goal is to predict the sentiment polarity $y_{i}$ and probability $P_{i}$ for the specific aspect $a_{i}$ conditioned on the $(v_{i},s_{i})$, i.e., $(y_{i},P_{i})=f(v_{i},s_{i},a_{i})$. For sentence-level MSA tasks, the $y_{i}$ and $P_{i}$ are predicted by sentence $s_{i}$ alongside the image $v_{i}$, i.e., $(y_{i},P_{i})=f(v_{i},s_{i})$.
General-Purpose Framework
To verify that our WisdoM works well on arbitrary architectures, we also apply WisdoM to LVLMs. We follow Wang et al. (2023) to construct the task instructions $I_{aspect}$ and $I_{sentence}$ for each task to elicit its ability to the corresponding task. The task instructions are presented as single-choice questions with well-formatted options (Shown in Fig. 2). The LVLMs can be seen as a sentiment classifier $f(\cdot)$. Therefore, the aspect-level task and sentence-level task can be formulated as $(y_{i},P_{i})=f(I_{aspect},v_{i},s_{i},a_{i})$ and $(y_{i},P_{i})=f(I_{sentence},v_{i},s_{i})$ respectively.
4 Methodology
Overview
Fig. 3 illustrates the overview of our method following three stages. In the ❶Prompt Templates Generation, we use ChatGPT to provide prompt templates. These prompt templates are fed into the LVLM with sentence $s_{i}$ and image $v_{i}$ to generate context, also called the ❷Context Generation. During ❸Contextual Fusion, we first compute the confidence, determining if the sample is uncertain (referred to as a hard sample). For hard samples, we fuse the predicted probability $P_{i}$ with $\hat{P}_{i}$ which obtained by incorporating context. Otherwise, we use $P_{i}$ as the final prediction.
4.1 Stage 1: Prompt Templates Generation
The main purpose of this stage is to design the prompt templates used to generate the context, so that the LVLM can better understand our intention and thus provide a more comprehensive contextual world knowledge. Inspired by Jiao et al. (2023); Zhong et al. (2023b), we ask ChatGPT to provide the appropriate prompt templates. The prompt templates provided by ChatGPT consider world knowledge of different perspectives, including history, social and cutural, etc. We insert a “Sentence: [x]” at the end of the prompt template to place the input sentence $s_{i}$. The example of prompt templates are shown in Appendix A.2.
4.2 Stage 2: Context Generation
In the context generation stage, prompt templates in ❶ are used to generate context which explicit world knowledge based on given the image $v_{i}$ and sentence $s_{i}$ by LVLMs Liu et al. (2023b); Ye et al. (2023). Specifically, we construct instruction by replacing the “[x]” in the prompt template with the sentence $s_{i}$. In addition, different LVLMs require a special token to indicate where the image $v_{i}$ is inserted. Taking LLaVA Liu et al. (2023b) as an example, we insert a special token “<image>” at the beginning of the instruction.
4.3 Stage 3: Contextual Fusion
We use the sentence-level task as an example. After obtained the context, we can intuitively use predicted sentiment polarity $\hat{y}_{i}$ obtained by incorporating context, i.e., $(\hat{y}_{i},\hat{P}_{i})=f(v_{i},s_{i},context)$. However, the context may contains irrelevant information that could disturb performance. Therefore, we first determine hard sample and then fuse the $P_{i}$ and $\hat{P}_{i}$ in hard sample.
Determine Hard Sample
Inspired by Zhang et al. (2021), we found that the ambiguous hard sample are commonly found around boundary areas of the sentiment polarity (e.g., the boundary areas of negative and neutral). Therefore, given a sample $m_{i}$, we only consider the difference $\delta_{i}$ between the highest and the second highest probabilities to determine whether is a hard sample:
$\delta_{i}=2\times max(P_{i})+min(P_{i})-1.$ (1)
Then, we denote uncertain threshold $\alpha$ to select samples that not exceed $\alpha$ as hard, i.e., $\mathcal{V}_{hard}=\{m_{i}|\delta_{i}\leq\alpha\}$, where we leave $\alpha=0.3$ as default.
Fusion with Context
Inspired by Li et al. (2022); O’Brien and Lewis (2023), we take the convex combinations of $P_{i}$ and $\hat{P}_{i}$ to obtain the final prediction $\tilde{P}_{i}$ for hard sample $m_{i}\in\mathcal{V}_{hard}$:
$\tilde{P}_{i}=P_{i}+\beta\cdot(\hat{P}_{i}-P_{i}),$ (2)
where $\beta$ is an interpolation coefficient. Intuitively, $(\hat{P}_{i}-P_{i})$ represent the information incorporating by context. $\beta$ is used to control the proportion of information introduced in context. When $\beta\rightarrow 0$, the effect brought by context is completely ignored and vice versa. Note that, we use $(y_{i},P_{i})$ as the final prediction when $m_{i}\notin\mathcal{V}_{hard}$.
5 Experiment
In this section, we apply WisdoM to aspect-level and sentence-level MSA tasks to verify its effectiveness and conduct extensive analysis to better understand the proposed method.
Method Twitter2015 Twitter2017
Acc. F1 Acc. F1
ESAFN Yu et al. (2019) 73.4 67.4 67.8 64.2
TomBERT Yu and Jiang (2019) 77.2 71.8 70.5 68.0
CapTrBERT Khan and Fu (2021) 78.0 73.2 72.3 70.2
JML Ju et al. (2021) 78.7 - 72.7 -
VLP-MABSA Ling et al. (2022) 78.6 73.8 73.8 71.8
CMMT Yang et al. (2022) 77.9 - 73.8 -
MMICL Zhao et al. (2023)* 76.0 72.7 74.1 74.0
-w/ WisdoM 77.3 (+1.3) 74.2 (+1.5) 75.7 (+1.6) 75.7 (+1.1)
[…]
-w/ WisdoM 78.9 (+1.0) 75.6 (+1.3) 75.6 (+1.3) 75.3 (+0.7)
[…]
-w/ WisdoM 77.3 (+0.5) 73.4 (+1.1) 74.5 (+0.3) 73.7 (+0.7)
[…]
-w/ WisdoM 81.5 (+1.5) 78.1 (+2.9) 77.6 (+1.7) 76.8 (+2.3)
Table 1: Comparison of our method (upon several advanced models) with existing works on Twitter2015 and 2017 benchmarks. The highest results are highlighted in bold, and * indicates the reproduced results.
Method MSED
Precision Recall Mac-F1
DCNN+AlexNet Jia et al. (2022) 71.02 70.09 70.31
DCNN+ResNet Jia et al. (2022) 74.73 74.73 74.64
BiLSTM+AlexNet Jia et al. (2022) 78.73 79.22 78.89
BERT+AlexNet Jia et al. (2022) 83.22 83.11 83.16
Multimodal Transformer Jia et al. (2022) 83.56 83.45 83.50
ALMT Zhang et al. (2023)* 83.73 83.98 83.73
-w/ WisdoM 89.62 (+5.89) 89.51 (+5.53) 89.54 (+5.81)
[…]
-w/ WisdoM 89.92 (+3.68) 88.24 (+1.69) 88.85 (+2.68)
[…]
-w/ WisdoM 88.72 (+0.97) 89.35 (+1.07) 88.95 (+0.97)
[…]
-w/ WisdoM 90.58 (+1.60) 90.41 (+1.64) 90.48 (+1.73)
Table 2: Performance of applying our WisdoM to advanced models on MSED benchmark, with reference results from existing works. The best results are bolded, and the * denotes the reproduced results.
5.1 Experimental Settings
Datasets
For aspect-level task, our two benchmark datasets are Twitter2015 and Twitter2017 Yu and Jiang (2019). Twitter2015 and Twitter2017 consist of multimodal tweets. Each multimodal tweet consists of text, an image posted alongside the tweet, aspects within the tweet, and the sentiment of each aspect. Each aspect is given a label from the set {negative, neutral, positive}. For sentence-level task, we evaluate our WisdoM on MSED Jia et al. (2022) dataset. MSED is a multimodal and multi-task dataset, which contains 9,190 text-image pairs.
Method Twitter2015 MSED
Acc. F1 Mac-F1
MMICL 76.0 72.7 86.2
 + context 69.5 72.1 86.8
 + Contextual Fusion 77.3 74.2 88.9
Table 3: Ablation study of context and its wise fusion module. We first only incorporate “context” and subsequently introduce the “contextual fusion” module.
Models
To demonstrate that our approach generalized across various architectures and sizes, we experimented on MMICL (14B) Zhao et al. (2023), LLaVA-v1.5 (13B) Liu et al. (2023a), mPLUG-Owl2 (8.2B) Ye et al. (2023), AoM (105M) Zhou et al. (2023), ALMT (112.5M) Zhang et al. (2023).
Evaluation Metrics
For aspect-level tasks, we use Accuracy (Acc) and F1 following previous studies Khan and Fu (2021); Ju et al. (2021); Ling et al. (2022); Zhou et al. (2023). For the sentence-level task, we adopt precision, recall, and macro-F1 (F1) as evaluation metrics.
Implement Details
For task-specific methods Zhou et al. (2023); Zhang et al. (2023), we employ the same inference procedures as their original works. For general-purpose methods Liu et al. (2023a); Zhao et al. (2023); Ye et al. (2023), we compute the likelihood that a LVLM generates the content of this choice given the question. We select the choice with the highest likelihood as model’s prediction. For fair comparison, these models are finetuned on target datasets. The learning rate is grid searched in {1e-5, 2e-5, 7.5e-5, 1e-4}, and batch size is in {16, 32, 256}. We optimize our model with AdamW Loshchilov and Hutter (2018).
Figure 4: Effect of interpolation coefficient $\beta$. We show how F1-score changes when varying the $\beta$ values.
Figure 5: Impact of uncertain threshold $\alpha$, illustrating how the F1-score changes when varying $\alpha$.
5.2 Hyperparameter Selection
In stage 3, Contextual fusion has two major hyperparameters $\beta$ and $\alpha$. To systematically study the impact of $\beta$ and $\alpha$, we first fix $\alpha=0.3$ and search different configurations for the $\beta$. Then we fix $\beta$ to the optimal value in the Twitter2015 dev set and search for $\alpha$. As shown in Fig. 4, we find that $\beta$ values within the range of $[0.4,0.5]$ demonstrate strong performance for both models. We observe that as the $\beta$ increases, indicating the introduction of more context information, all models exhibit a substantial decrease. We hypothesize that this decline may be caused by introducing excessive noise within the context. Fig. 5 shows the trend on Twitter2015 dev set for different $\alpha$ values. We note that performance remains strong for both models when $\alpha$ values are in the range of $[0.3,0.4]$. In addition, we observe no notable decrease in trends as $\alpha$ increases. One possible explanation for this phenomenon is that the introduced context proves to be effective for samples with low confidence, but its impact diminishes for samples with high confidence. For a fairer comparison, we fixed $\alpha=0.3$, and $\beta$ is the optimal value on the dev set for testing.
5.3 Main Results
Evaluation for Aspect-level MSA Task
We compare against SOTA on Twitter2015 and 2017. Table 1 shows the performance of aspect-level tasks. Our WisdoM achieves consistent and significant improvement on four models across two datasets. The WisdoM brings max 2.9% and 2.3% F1-gains on Twitter2015 and Twitter2017 respectively. The results show that our method has a clear advantage.
Evaluation for Sentence-level MSA Task
As shown in Table 2, the best F1 on MSED, our method WisdoM with LLaVA-v1.5 is 90.48, outperforming LLaVA-v1.5 (88.75), mPLUG-Owl2 (87.98), MMICL (86.17) and ALMT (83.73). The significant improvement on the SOTA MSA method ALMT brings 5.81% F1-gains. In addition, we observe a more pronounced enhancement in the effect on small-scale models, whereas the impact is less prominent on larger models. This suggests that contextual world knowledge is particularly crucial for small models in MSA.
5.4 Ablation Study
Effect of Different Modules.
To better understand the role of each module in our method WisdoM, Table 3 presents the ablation results of the gradual addition of different components. Compared with the baseline (MMICL), adding context results in performance degradation. We found that context containing irrelevant information leads to degradation of performance on Twitter2015. The further addition of Contextual Fusion brings a significant improvement of 2.1 F1 score.
Effect of Different Fusion Strategies.
In Table. 5.4, we explore different fusion strategies experiments. We conducted experiments involving $mean(P_{i},\hat{P}_{i})$ and $max(P_{i},\hat{P}_{i})$. We observe that we can achieve a marginal improvement with these straightforward fusion strategies. Furthermore, we calculate the Jensen-Shannon (JS) divergence of $P_{i}$ with the uniform distribution to serve as the fusion weight, i.e., $\beta$. However, JS divergence is influenced by the class with the smallest probability, introducing an impractical level of difficulty.
Table 4: Ablation study of different fusion strategies. “JS” denotes Jensen-Shannon divergence. “CF” denotes our Contextual Fusion.
Method Twitter2015
Acc. F1
AoM 80.00 75.20
 -w/ Average 80.33 (+0.33) 70.69 (-4.51)
 -w/ Max 80.56 (+0.56) 76.81 (+1.61)
 -w/ JS 80.78 (+0.78) 77.39 (+2.19)
 -w/ CF 81.45 (+1.45) 78.12 (+2.92)
[…]
 -w/ Average 75.53 (-0.45) 72.33 (-0.38)
 -w/ Max 75.42 (-0.56) 72.16 (-0.55)
 -w/ JS 77.09 (+1.11) 73.92 (+1.21)
 -w/ CF 77.32 (+1.34) 74.18 (+1.47)
Figure 6: Performance of scaling WisdoM on Twitter2015 with different a) model and b) data scales.
Table 4: Ablation study of different fusion strategies. “JS” denotes Jensen-Shannon divergence. “CF” denotes our Contextual Fusion.